OpenStudy (anonymous):

A point Q on a segment with endpoints A (2, −1) and C (4, 2) partitions the segment in a 3:1 ratio. Find Q. @jigglypuff314

jigglypuff314 (jigglypuff314):

ehhh *tilts head* 3:1 means you can split it into 4 parts graph the given points

OpenStudy (anonymous):

I dont understanf...

OpenStudy (anonymous):

:(

OpenStudy (bradely):

Q=((mx2+nx1)/(m+n),(my2+ny1)/(m+n) m=3,n=1 Source: http://www.mathskey.com/question2answer/

jigglypuff314 (jigglypuff314):

when you graph the points, you can see that there is a difference of 2 in the x aspect so 2 divided by 4 parts = 0.5 

OpenStudy (bradely):

A(x1,y1) C(x2,y2) x1=2,y1=-1 x2=4,y2=2

OpenStudy (anonymous):

uhm... and would 2, 2 come to be Q? @bradely

jigglypuff314 (jigglypuff314):

oh... math xD thanks @bradely that works

OpenStudy (bradely):

x= (3*4+1*2)/(3+1)=14/4 =7/2 x= (3*2+1*-1)/(3+1)=5/4

OpenStudy (anonymous):

@bradely now I'm confused... can you explain please

jigglypuff314 (jigglypuff314):

it is a formula he is using and just plugging in numbers with, which works :P

OpenStudy (bradely):

Points are A (2, −1) and C (4, 2) x1, y1 x2, y2 ratio m:n =3:1 Substitute the formula Q (x,y) = ((mx2+nx1)/(m+n),(my2+ny1)/(m+n)

OpenStudy (anonymous):

and what would Q be then??? I'm sry, i dont understand @bradely

OpenStudy (bradely):

Q is the dividing point of A and C Q has x and y coordinates x is (mx2+nx1)/(m+n) y is (my2+ny1)/(m+n)

OpenStudy (anonymous):

oh okay @bradely thank you, now i understand